31. Magpie Corporation uses the total cost concept of product pricing. Below is the cost information for the production...

31.

Magpie Corporation uses the total cost concept of product pricing. Below is the cost information for the production and sale of 56,300 units of its sole product. Magpie desires a profit equal to a 21% rate of return on invested assets of $643,000.

Fixed factory overhead cost $36,000
Fixed selling and administrative costs 7,900
Variable direct materials cost per unit 5.06
Variable direct labor cost per unit 1.88
Variable factory overhead cost per unit 1.13
Variable selling and administrative cost per unit 4.50

The MARKUP PERCENTAGE on total cost for the company's product is

a.30.5%

b.14.4%

c.18.0%

d.23.4%
Answer #1
Variable direct materials cost per unit 5.06
Variable direct labor cost per unit 1.88
Variable factory overhead cost per unit 1.13
Fixed factory overhead cost per unit 0.64 =36000/56300
Variable selling and administrative cost per unit 4.50
Fixed selling and administrative cost per unit 0.14 =7900/56300
Total cost per unit 13.35
Desired Profit per unit 2.40 =(643000*21%)/56300
Desired Profit per unit 2.40
Divide by Total cost per unit 13.35
Markup Percentage on Total cost 18.0%
Option C 18.0% is correct